Suppose an ambulance is driving at 28m/s, with its siren going off at a frequency of 700Hz. You are in your car, waiting at a red light as it approaches. Assume the sound of the siren is traveling at 341m/s.

a. What frequency do you hear the horn as it gets closer to you?

b. What frequency do you hear the horn as it passes you?
Answer #1

Here, we can use Doppler effect

________________

(a)

we will hear higher frequency

f = fo ( v + vobserver / v - vsource )

f = 700 ( 341 + 0 / 341 - 28)

f = 762.62 H

__________________

(b)

we will hear lower frequency

f = fo ( v + vobserver / v + vsource )

f = 700 ( 341 + 0 / 341 + 28)

f = 646.9 Hz